### 15 Dec 2023 v2

@@ Abstract
Motivation: Benefiting from the advancements in deep learning, various genomic analytical techniques, such as survival analysis, classification of tumors and their subtypes, and exploration of specific pathways, have significantly enhanced our understanding of the biological mechanisms driving cancer. However, the overfitting issue, arising from the limited number of patient samples, poses a challenge in improving the accuracy of genome analysis by deepening the neural network. Furthermore, it remains uncertain whether novel approaches such as the sparsely gated mixture of expert (MOE) and self-attention mechanisms can improve the accuracy of genomic analysis. In this paper, we introduce a novel sparsely gated RNA-seq analysis framework called Gene-MOE. This framework exploits the potential of the MOE layers and the proposed mixture of attention expert (MOAE) layers to enhance the analysis accuracy. Additionally, it addresses overfitting challenges by integrating pan-cancer information from 33 distinct cancer types through pre-training.
Results: We pre-trained Gene-MOE on TCGA pan-cancer RNA-seq dataset with 33 cancer types. Subsequently, we conducted experiments involving cancer classification and survival analysis based on the pre-trained Gene-MOE. According to the survival analysis results on 14 cancer types, Gene-MOE outperformed state-of-the-art models on 12 cancer types. Through detailed feature analysis, we found that the Gene-MOE model could learn rich feature representations of high-dimensional genes. According to the classification results, the total accuracy of the classification model for 33 cancer classifications reached 95.8%, representing the best performance compared to state-of-the-art models. These results indicate that Gene-MOE holds strong potential for use in cancer classification and survival analysis.
Availability and implementation: Our implementation of Gene-MOE is available at https://github.com/menggerShe
rry/gene-moe.
1 Introduction
@@
In the past two decades, a series of machine learning methods have been proposed for genomic analysis. Representative work includes gene survival analysis using the COX regression model (Lin and Wei, 1989), cancer classification and tumor biomarker identification using random forest (Capper et al., 2018; Díaz-Uriarte and Alvarez de Andrés, 2006), cluster gene expression data using principal component analysis (PCA) (Yeung and Ruzzo, 2001), gene selection and cancer classification using support vector machine (SVM) (Duan et al., 2005), and gene expression pattern classification using the linear regression method (Liu et al., 2019). Owing to their simplicity and user-friendliness, machine learning models have proved to be highly efficient in processing genomic analysis tasks, delivering commendable results across various functions, such as dimensional reduction and clustering visualization. However, the nonlinear biological characteristics inherent in high-dimensional genes cannot be effectively learned by the machine learning method, leading to precision loss in genomic analysis.
@@
In recent years, deep learning methods have received widespread attention. With powerful fitting capabilities and more accurate prediction effects than traditional learning methods, deep learning methods can more effectively handle the potential correlation features of high-dimensional data. Many researchers have adopted deep learning methods for feature extraction and downstream prediction. Representative work includes the Cox-nnet model (Ching et al., 2018), DeepCC (Gao et al., 2019), DeepCues (Zeng et al., 2021), PathCNN (Oh et al., 2021), and the cancer prognosis and classification method using a graph convolutional network (GCN) (Ramirez et al., 2020, 2021). Unlike machine learning methods, deep learning methods have greatly improved fitting performance and prediction accuracy, thus improving cancer diagnosis and prognosis. However, owing to the imbalance of samples between different cancer types and the few fitting samples for cancer types such as uterine carcinosarcoma (UCS), ocular melanoma (UVM), and large b-cell lymphoma (DLBC), deep learning methods must build two to four layers of a deep neural network (DNN) to prevent over-fitting problems, so the model cannot learn the deep correlations between high-dimensional genes.
@@
Pre-trained models have received widespread attention owing to their powerful feature learning capabilities. These models require training a network with billions of parameters to learn common features across various data samples and transferring weights to a specific task to capture the unique properties. Some notable models are transformer-based models (Li et al., 2023; Wang et al., 2022a, b), which were first proposed in the natural language processing field. They include dense models such as GPT-3 (Brown et al., 2020), Gopher (Rae et al., 2021), and the sparse model based on the mixture of expert (MOE) model (Shazeer et al., 2017; DBLP:conf/iclr/LepikhinLXCFHKS31). Moreover, many pre-training models have been introduced to solve biological problems and have achieved remarkable results, including BioBERT (Lee et al., 2020), DNABERT (Ji et al., 2021), and scBERT (Yang et al., 2022). However, challenges persist in the application of pre-trained models to high-dimensional genetic data. First, the number of genes far exceeds the number of patient samples, and most genes contain no useful information for diagnosis and prognosis, exacerbating the risk of over-fitting during training. The TCGA dataset is estimated to contain over 50,000 genes, of which protein-coding genes make up approximately 21,000 (Pertea et al., 2018), and most genes contain no useful information. Appropriate data preprocessing and augmentation methods must be designed to avoid meaningless information as much as possible. Second, a variety of genes often express together to cause cancer, and a strong correlation exists between different genes. Exploring such correlations from high-dimensional genetic features requires special consideration to design appropriate feature extraction methods tailored to genomic data.
@@
The rest of this paper is organized as follows. Section 2 illustrates our method, including the dataset preparation, Gene-MOE framework illustration, and training strategy. Section 3 presents our experimental results for Gene-MOE. Finally, in Section 4, we provide our conclusion.
2 Methods
2.1 Dataset Preparation
@@
In this study, we primarily utilized the pan-cancer RNA-seq database from the TCGA dataset of the Pan-Cancer Atlas project, which consists of 33 cancer types. Furthermore, we downloaded the specific RNA-seq dataset for each TCGA cancer type. UCSC Xena already includes the preliminary processing for these 34 datasets, so we directly used the RNA-seq data after the FPKM-UQ normalization from UCSC Xena. In the initial dataset, each patient was associated with 60,484 genes, a majority of which exhibited null expression and had no relevance to cancer. Consequently, our first step involved filtering out these non-contributing genes. We first deleted the empty genes and selected overlapping genes among 34 datasets. Next, we filtered out genes with variances less than 0.4 and mean values less than 0.8 according to (Ramirez et al., 2021). The reason is that genes that meet this low variance and low mean range were 0 in most patients and did not make any contribution to model training. Finally, before feeding the data into the network, we performed min-max normalization. After preparation, we selected 25,182 genes for each dataset. Figure 1 shows the patient number of these 33 cancer types.
2.2 Framework Illustration
@@
2.2.1 Sparsely Gated MOE Layer
@@
2.2.2 Mixture of Attention Expert Layer
@@ 2.3 Training Strategy
2.3.1 Data Augmentation by Gaussian Noise
@@
2.3.2 Joint Training Using a Generative Adversarial Network
@@ 2.3.3 Measuring the Distribution
2.3.4 Measuring the Similarity of Genes
@@
2.3.5 Balancing Expert Utilization
@@
where $B$ denotes the batch size of the input features, and $G$ denotes the gating layer. To further improve the balanced loading, we also introduced the load balance loss $\mathcal{L}_{load}$ in (Shazeer et al., 2017).
2.3.6 Overall Pre-training Loss
@@ 2.4 Experimental Settings
After the pre-training phase, we combined the backbone of the Gene-MOE with the classifier to perform survival analysis and cancer classification tasks. During the survival analysis phase, we evaluated the survival model in the same way as Cox-nnet (Ching et al., 2018) and the VAE-Cox (Kim et al., 2020). We trained specific Gene-MOE on 14 TCGA datasets of common cancer types. For each dataset, we used the same way to randomly divide the train dataset and test dataset as a ratio of 4:1. The optimal hyperparameters of each model were selected using the Bayesian Optimization strategy. Moreover, we repeated the entire process 5 times and calculated the average result to avoid the bias of the random splitting. In the cancer classification phase, due to the imbalance of samples for each cancer, directly dividing the pan-cancer data into the training set and the test set at a ratio of 4:1 will result in the missing cancer types with a small number of samples in the test set. To solve the above issues, we introduced a new partition method that the sample set of each cancer was divided into the train subset and test subset according to the ratio of 4:1, and then the training subsets and test subsets of 33 cancers were combined to obtain the training set and test set of the classification task. We also used the Bayesian Optimization strategy to search the best hyperparameters and repeated the entire classification process 5 times to calculate the average result.
@@
2.5 Evaluation Metric
In the survival analysis phase, the evaluation method we mainly used was the Concordance Index (Steck et al., 2007), which is widely used in survival analysis models and ranges from 0 to 1. When the Concordance Index $\leq$ 0.5, the model has completed an ineffective survival analysis prediction. When the Concordance index $\geq$ 0.5 and higher, the prediction effect of the model has been better. In the cancer classification phase, we mainly used the accuracy, precision, recall, and F1-score. The accuracy is denoted as
@@ 2.5 Evaluation Metric
3 Results
3.1 Performance on Survival Analysis
@@
We selected 14 representative cancer types and analyzed the performance of our method in comparison with two state-of-the-art models: Cox-nnet (Ching et al., 2018) and VAE-Cox (Kim et al., 2020). Figure 5 shows the Concordance Index on 14 cancer types. Compared with Cox-nnet and the VAECox, the Gene-MOE outperformed on 12 cancer types, with a higher mean Concordance Index, which shows that Gene-MOE carried out a more accurate survival analysis compared with these two models.
We further employed Gene-MOE for survival analysis predictions. Specifically, we selected the test datasets of the 12 cancer types indicated in Figure 5. Patients were divided into high-risk and low-risk groups based on the average predictions of Gene-MOE. Subsequently, we plotted Kaplan–Meier (KM) survival curves and conducted log-rank tests. We also carried out the same survival analysis for VAE-Cox. Figure S1 displays the KM curve results for 12 cancer types. We found that Gene-MOE outperformed VAE-Cox significantly for these 12 cancer types with a lower logP value, indicating the effective capability of Gene-MOE to split patients into high-risk and low-risk groups.
3.2 Performance on Cancer Type Classification
We assessed the performance of Gene-MOE through a cancer classification task. Specifically, we designed a new classification head that takes the Gene-MOE’s backbone network as input, predicting probabilities for 33 distinct cancer types. Then Focal loss (Lin et al., 2017) was employed during the training process to mitigate the imbalance issue and enhance the model performance. Subsequently, we evaluated the precision, recall, and F1-score of the classification model employing Gene-MOE. The performance results are depicted in Figure 4. As illustrated in this figure, our model demonstrated significant performance on 32 cancer types except rectum adenocarcinoma (READ), with a total accuracy of 95.8%.
To prove the novelty of proposed classifier, we selected several state-of-the-art classifiers for comparison, including classifier models based on machine learning methods such as random forest (Statnikov et al., 2008) and SVM (Kourou et al., 2015; Statnikov et al., 2008), as well as deep learning based classifier models like MLP (Ahn et al., 2018) and CNN (Mostavi et al., 2020). To ensure the fairness of the comparative experiments, we conducted five independent processes and calculated the average of the classification results to avoid the bias. Table 1 shows the classification result of these five models. According to table 1, We found that the classification results of the Gene-MOE model on the test dataset was improved by 0.1-0.5 compared with the other four methods.
Table 1: Comparison of classification metrics
[…]
@@ 3.2 Performance on Cancer Type Classification
We further analyzed the classification performance by constructing a confusion matrix. Figure S2 illustrates the confusion matrix constructed for 33 types of cancer. Our model demonstrated accurate classification of all 33 cancer types with a notably low misclassification rate according to the confusion matrix. Moreover, we observed that our model misclassified 29 cases of READ as colon adenocarcinoma (COAD). This issue occurred because READ and COAD are genetically identical (Network et al., 2012), and COAD samples outnumbered READ samples, causing the model to misclassify some READ patients as COAD patients.
3.3 Feature Analysis of Gene-MOE
3.3.1 Correlation Analysis
@@
We conducted a correlation analysis between the low-dimensional features extracted by the Gene-MOE model and the high-dimensional genes of patients. Specifically, we used 1,000 patient samples from the test dataset to feed into the Gene-MOE model, and the model evaluated the low-dimensional features of each patient. We then selected the top 20 features with the highest variance as the leading features and calculated their Pearson correlations with the original patient genetic information. The results are shown in Figure S3(a), which presents a heat map illustrating the correlation between these leading features and the genes of 1,000 patients, which unequivocally indicates a significant correlation between the low-dimensional features predicted by the Gene-MOE model and the original input features. Based on Figure S3(a), we further refined our analysis by identifying genes with an average absolute correlation exceeding 0.4 with respect to the 20 leading features. Then, 29 strongly correlated genes were filtered, which is shown in Figure S3(b). According to the result in Figure S3(b), we observed that many genes with strong correlations to the leading features were cancer-related genes.
3.3.2 Visualization Analysis

### 18 Dec 2023 v3

@@ Abstract
Deep learning based genomic analysis methods enhanced our understanding for cancer research. However, the overfitting issue, arising from the limited number of patient samples, presents a challenge in improving the accuracy of genome analysis by deepening the neural network. Furthermore, it remains uncertain whether novel approaches such as the sparsely gated mixture of expert (MOE) and self-attention mechanisms can improve the accuracy of genomic analysis. We introduce a novel sparsely gated prognosis and classification analysis framework called Gene-MOE. This framework exploits the potential of the MOE layers and the proposed mixture of attention expert (MOAE) layers to enhance the analysis accuracy. Additionally, it addresses overfitting challenges by integrating pan-cancer information from 33 distinct cancer types through pre-training. According to the survival analysis results on 14 cancer types, Gene-MOE outperformed state-of-the-art models on 12 cancer types. Through detailed feature analysis, we found that the Gene-MOE model could learn rich feature representations of high-dimensional genes. According to the classification results, the total accuracy of the classification model for 33 cancer classifications reached 95.8%, representing the best performance compared to state-of-the-art models. These results indicate that Gene-MOE holds strong potential for use in cancer classification and survival analysis.
Index Terms:
Survival analysis, Cancer classification, Genomic analysis, Mixture of expert, Self-attention.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Carcinogenic manifestations often result from mutations in one or more genes [1]. The Cancer Genome Atlas (TCGA) project represents a major advance in cancer genomics. The tens of thousands of pretreatment samples in TCGA, including over 30 cancer types and numerous measurements, including RNA sequencing (RNA-seq), DNA methylation, and copy-number variation, deepen our understanding of cancer-related genes and their clinical relevance [2, 3, 4]. To find more meaningful biological functions, researchers have developed methods such as prognosis prediction [5, 6], tumor subtypes [7], microsatellite instability (MSI) [8], immunological aspects [9], and certain pathways of interest [10].
@@ I Introduction
In the past two decades, a series of machine learning methods have been proposed for genomic analysis. Representative work includes gene survival analysis using the COX regression model [11], cancer classification and tumor biomarker identification using random forest [12, 13], cluster gene expression data using principal component analysis (PCA) [14], gene selection and cancer classification using support vector machine (SVM) [15], and gene expression pattern classification using the linear regression method [16]. Owing to their simplicity and user-friendliness, machine learning models have proved to be highly efficient in processing genomic analysis tasks, delivering commendable results across various functions, such as dimensional reduction and clustering visualization. However, the nonlinear biological characteristics inherent in high-dimensional genes cannot be effectively learned by the machine learning method, leading to precision loss in genomic analysis.
@@
In recent years, deep learning methods have received widespread attention. With powerful fitting capabilities and more accurate prediction effects than traditional learning methods, deep learning methods can more effectively handle the potential correlation features of high-dimensional data. Many researchers have adopted deep learning methods for feature extraction and downstream prediction. Representative work includes the Cox-nnet model [17], DeepCC [18], DeepCues [19], PathCNN [20], and the cancer prognosis and classification method using a graph convolutional network (GCN) [21, 22]. Unlike machine learning methods, deep learning methods have greatly improved fitting performance and prediction accuracy, thus improving cancer diagnosis and prognosis. However, owing to the imbalance of samples between different cancer types and the few fitting samples for cancer types such as uterine carcinosarcoma (UCS), ocular melanoma (UVM), and large b-cell lymphoma (DLBC), deep learning methods must build two to four layers of a deep neural network (DNN) to prevent over-fitting problems, so the model cannot learn the deep correlations between high-dimensional genes. Recently, the pre-trained models have received widespread attention owing to their powerful feature learning capabilities. These models require training a network with billions of parameters to learn common features across various data samples and transferring weights to a specific task to capture the unique properties. Some notable models are transformer-based models [23, 24], which were first proposed in the natural language processing field. They include dense models such as GPT-3 [25], Gopher [26], and the sparse model based on the mixture of expert (MOE) model [27, 28]. Moreover, many pre-training models have been introduced to solve biological problems and have achieved remarkable results, including BioBERT [29], DNABERT [30], and scBERT [31].
@@
However, challenges persist in the application of pre-trained models to high-dimensional genetic data. First, the number of genes far exceeds the number of patient samples, and most genes contain no useful information for diagnosis and prognosis, exacerbating the risk of over-fitting during training. The TCGA dataset is estimated to contain over 50,000 genes, of which protein-coding genes make up approximately 21,000 [32], and most genes contain no useful information. Appropriate data preprocessing and augmentation methods must be designed to avoid meaningless information as much as possible. Second, a variety of genes often express together to cause cancer, and a strong correlation exists between different genes. Exploring such correlations from high-dimensional genetic features requires special consideration to design appropriate feature extraction methods tailored to genomic data.
@@
The rest of this paper is organized as follows. Section II illustrates our method, including the dataset preparation, Gene-MOE framework illustration, and training strategy. Section III presents our experimental results for Gene-MOE. Finally, in Section IV, we provide our conclusion.
II Methods
II-A Dataset Preparation
@@
In this study, we primarily utilized the pan-cancer RNA-seq database from the TCGA dataset of the Pan-Cancer Atlas project, which consists of 33 cancer types. Furthermore, we downloaded the specific RNA-seq dataset for each TCGA cancer type. UCSC Xena already includes the preliminary processing for these 34 datasets, so we directly used the RNA-seq data after the FPKM-UQ normalization from UCSC Xena. In the initial dataset, each patient was associated with 60,484 genes, a majority of which exhibited null expression and had no relevance to cancer. Consequently, our first step involved filtering out these non-contributing genes. We first deleted the empty genes and selected overlapping genes among 34 datasets. Next, we filtered out genes with variances less than 0.4 and mean values less than 0.8 according to [22]. The reason is that genes that meet this low variance and low mean range were 0 in most patients and did not make any contribution to model training. Finally, before feeding the data into the network, we performed min-max normalization. After preparation, we selected 25,182 genes for each dataset. Figure 1 shows the patient number of these 33 cancer types.
II-B Framework Illustration
@@
II-B1 Sparsely Gated MOE Layer
@@
II-B2 Mixture of Attention Expert Layer
@@ II-C Training Strategy
II-C1 Data Augmentation by Gaussian Noise
@@
II-C2 Joint Training Using a Generative Adversarial Network
@@ II-C3 Measuring the Distribution
II-C4 Measuring the Similarity of Genes
@@
II-C5 Balancing Expert Utilization
@@
where $B$ denotes the batch size of the input features, and $G$ denotes the gating layer. To further improve the balanced loading, we also introduced the load balance loss $\mathcal{L}_{load}$ in [27].
II-C6 Overall Pre-training Loss
@@ II-D Experimental Settings
After the pre-training phase, we combined the backbone of the Gene-MOE with the classifier to perform survival analysis and cancer classification tasks. During the survival analysis phase, we evaluated the survival model in the same way as Cox-nnet [17] and the VAE-Cox [33]. We trained specific Gene-MOE on 14 TCGA datasets of common cancer types. For each dataset, we used the same way to randomly divide the train dataset and test dataset as a ratio of 4:1. The optimal hyperparameters of each model were selected using the Bayesian Optimization strategy. Moreover, we repeated the entire process 5 times and calculated the average result to avoid the bias of the random splitting. In the cancer classification phase, due to the imbalance of samples for each cancer, directly dividing the pan-cancer data into the training set and the test set at a ratio of 4:1 will result in the missing cancer types with a small number of samples in the test set. To solve the above issues, we introduced a new partition method that the sample set of each cancer was divided into the train subset and test subset according to the ratio of 4:1, and then the training subsets and test subsets of 33 cancers were combined to obtain the training set and test set of the classification task. We also used the Bayesian Optimization strategy to search the best hyperparameters and repeated the entire classification process 5 times to calculate the average result.
@@
II-E Evaluation Metric
In the survival analysis phase, the evaluation method we mainly used was the Concordance Index [34], which is widely used in survival analysis models and ranges from 0 to 1. When the Concordance Index $\leq$ 0.5, the model has completed an ineffective survival analysis prediction. When the Concordance index $\geq$ 0.5 and higher, the prediction effect of the model has been better. In the cancer classification phase, we mainly used the accuracy, precision, recall, and F1-score. The accuracy is denoted as
@@ II-E Evaluation Metric
III Results
III-A Performance on Survival Analysis
@@
Figure 6: Kaplan–Meier survival curves using MOE-Cox and VAE-cox on 12 cancer types. The Gene-MOE model shows a lower logP value on 10 cancer data sets. It shows that the Gene-MOE can predict hazard ratios that divide into high-risk groups and low-risk groups more significantly compared with VAE-cox on 10 cancer datasets.
We selected 14 representative cancer types and analyzed the performance of our method in comparison with two state-of-the-art models: Cox-nnet [17] and VAE-Cox [33]. Figure 5 shows the Concordance Index on 14 cancer types. Compared with Cox-nnet and the VAECox, the Gene-MOE outperformed on 12 cancer types, with a higher mean Concordance Index, which shows that Gene-MOE carried out a more accurate survival analysis compared with these two models.
We further employed Gene-MOE for survival analysis predictions. Specifically, we selected the test datasets of the 12 cancer types indicated in Figure 5. Patients were divided into high-risk and low-risk groups based on the average predictions of Gene-MOE. Subsequently, we plotted Kaplan–Meier (KM) survival curves and conducted log-rank tests. We also carried out the same survival analysis for VAE-Cox. Figure 6 displays the KM curve results for 12 cancer types. We found that Gene-MOE outperformed VAE-Cox significantly for these 12 cancer types with a lower logP value, indicating the effective capability of Gene-MOE to split patients into high-risk and low-risk groups.
III-B Performance on Cancer Type Classification
We assessed the performance of Gene-MOE through a cancer classification task. Specifically, we designed a new classification head that takes the Gene-MOE’s backbone network as input, predicting probabilities for 33 distinct cancer types. Then Focal loss [35] was employed during the training process to mitigate the imbalance issue and enhance the model performance. Subsequently, we evaluated the precision, recall, and F1-score of the classification model employing Gene-MOE. The performance results are depicted in Figure 4. As illustrated in this figure, our model demonstrated significant performance on 32 cancer types except rectum adenocarcinoma (READ), with a total accuracy of 95.8%.
To prove the novelty of proposed classifier, we selected several state-of-the-art classifiers for comparison, including classifier models based on machine learning methods such as random forest [36] and SVM [6, 36], as well as deep learning based classifier models like MLP [37] and CNN [38]. To ensure the fairness of the comparative experiments, we conducted five independent processes and calculated the average of the classification results to avoid the bias. Table I shows the classification result of these five models. According to table I, We found that the classification results of the Gene-MOE model on the test dataset was improved by 0.1-0.5 compared with the other four methods.
TABLE I: Comparison of classification metrics
Accuracy Precision Recall F1-Score
RandomForest 0.9010 0.8943 0.8338 0.8409
@@ III-B Performance on Cancer Type Classification
We further analyzed the classification performance by constructing a confusion matrix. Figure 7 illustrates the confusion matrix constructed for 33 types of cancer. Our model demonstrated accurate classification of all 33 cancer types with a notably low misclassification rate according to the confusion matrix. Moreover, we observed that our model misclassified 24 cases of READ as colon adenocarcinoma (COAD). This issue occurred because READ and COAD are genetically identical [39], and because COAD samples outnumbered READ samples, causing the model to misclassify some READ patients as COAD patients.
Figure 7: Confusion matrix of test samples predicted by classification model using Gene-MOE with 33 cancer types.
Figure 8: Feature analysis of Gene-MOE. (a) Pearson correlation heat map between the leading feature and the patient genes. The features learned by Gene-MOE were strongly correlated with patient genes. (b) Pearson correlation heat map by selecting mean absolute coefficient greater than $0.5$. (c) TSNE and UMAP results of real genes and Gene-MOE reconstructed genes. According to the results, the reconstructed genes maintained the same distribution as the real genes.
III-C Feature Analysis of Gene-MOE
III-C1 Correlation Analysis
@@
We conducted a correlation analysis between the low-dimensional features extracted by the Gene-MOE model and the high-dimensional genes of patients. Specifically, we used 1,000 patient samples from the test dataset to feed into the Gene-MOE model, and the model evaluated the low-dimensional features of each patient. We then selected the top 20 features with the highest variance as the leading features and calculated their Pearson correlations with the original patient genetic information. The results are shown in Figure 8(a), which presents a heat map illustrating the correlation between these leading features and the genes of 1,000 patients, which unequivocally indicates a significant correlation between the low-dimensional features predicted by the Gene-MOE model and the original input features. Based on Figure 8(a), we further refined our analysis by identifying genes with an average absolute correlation exceeding 0.4 with respect to the 20 leading features. Then, 29 strongly correlated genes were filtered, which is shown in Figure 8(b). According to the result in Figure 8(b), we observed that many genes with strong correlations to the leading features were cancer-related genes. For example, the TMPRSS4 gene is an emerging potential therapeutic target in cancer  [40]. Moreover, tensin4 expression showed prognostic relevance in gastric cancer [41]. Furthermore, E2F1-initiated transcription of PRSS22 promoted breast cancer metastasis by cleaving ANXA1 and activating the FPR2/ERK signaling pathway [42]. In addition, down-regulation of FXYD3 expression was observed in lung cancers [43]. Moreover, ST14 gene expression affected breast cancer [44, 45, 46]. In addition, RAB25 has been implicated in various cancers, with reports presenting it as both an oncogene and a tumor-suppressor gene [47, 48]. Long intergenic non-coding RNA 00324 promoted gastric cancer cell proliferation by binding with HuR and stabilizing FAM83B expression [49]. CBLC expression was found to be higher in breast cancer tissues and cells than in normal tissues and cells [50]. Furthermore, Serpin B5 was shown to be a CEA‐interacting biomarker for colorectal cancer [51]. PKP3 interactions with the MAPK-JNK-ERK1/2-mTOR pathway regulated autophagy and invasion in ovarian cancer [52]. In addition, LAD1 expression was associated with the metastatic potential of colorectal cancer cells [53]. ELF3 was found to be a negative regulator of epithelial–mesenchymal transition in ovarian cancer cells [54]. Moreover, the expression patterns of CEACAM5 and CEACAM6 were observed in primary and metastatic cancers [55]. Grhl2 determined the epithelial phenotype of breast cancers and promoted tumor progression [56]. KLF5 promoted breast cancer proliferation, migration, and invasion, in part by up-regulating the transcription of TNFAIP2 [57]. Finally, genetic predisposition to colon and rectal adenocarcinoma was found to be mediated by a super-enhancer polymorphism coactivating CD9 and PLEKHG6 [58].
III-C2 Visualization Analysis